OpenStudy (anonymous):

The perimter of the rectangle is thirty feet. The length is three feet longer than the width. Find the demensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. What is length? What is width?

OpenStudy (anonymous):

Okay so lets draw a picture, pictures always help in questions like this. So tell me what perimeter is..

OpenStudy (anonymous):

i have 2(x+y)=thirty x= y+three 2(y+three+y)=thirty 2y+three=27

OpenStudy (anonymous):

i think its 2 by 25

OpenStudy (anonymous):

I got a different answer, so this is what I did:  So y=x+3 and 2(y+x)=30 (the perimeter) We can substitute y=x+3 into the second equation so it will be 2y+2x=30 so 2(x+3)+2x=30 which is 2x+6+2x=30 so 4x+6=30 subtract 6 4x=24 divide 24 by 4 x=6

OpenStudy (anonymous):

You can then plug x back into the equation to solve for Y

OpenStudy (anonymous):

ok so the lentgh is 24 the width is 2

OpenStudy (anonymous):

i dont know how i got 25

OpenStudy (anonymous):

i think i miss count

OpenStudy (anonymous):

How'd you get 2?

OpenStudy (anonymous):

for the width is x 2 or well idk

OpenStudy (anonymous):

if x is the width and x=6 then 6 is the width If the length is the width + 3 then x+3=y (y is the length) 6+3=9 so the width is 6 and the length is 9? confused by that?

OpenStudy (anonymous):

ok i seee 4/24 is 6

OpenStudy (anonymous):

nope i see my errors but i think i was getting warm lol

OpenStudy (anonymous):

and if you add the sides, 6+6+9+9 that equals 30, which is what the question said so you can check that way.

OpenStudy (anonymous):

you add the three feet to 6 to get 9

OpenStudy (anonymous):

Yup, since the problem said "The length is three feet longer than the width."

OpenStudy (anonymous):

ok it says the first equation 2x+2y=? and second equat x=y+?

OpenStudy (anonymous):

so i put thirty and three for the 1st and 2nd equat where the ? marks are at...

OpenStudy (anonymous):

2x+2y=thirty x=y+three

OpenStudy (anonymous):

do you see?

OpenStudy (anonymous):

It should be y=x+3 if you're doing by my diagram, but you just switched the x's and y's around. if you solve what you just put you should get the same answers

OpenStudy (anonymous):

oh ok thats ok it still camee out right

OpenStudy (anonymous):

So you know what the width and length are?

OpenStudy (anonymous):

6 and 9 right

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

yeeeaaa

OpenStudy (anonymous):

you are smart i am trying at last i attempt to try do you havee time for 2 more prob?

OpenStudy (anonymous):

Sure! I'm not sure if you should do this but clicking best repose may show others where the answer is encase they're looking at this problem. but Yeah I have time for some more.

OpenStudy (anonymous):

ok i didnt know ok the probelm is Shade the solution set to th system of inequality. x+three y>-2 2x+ ygreater than = to three

OpenStudy (anonymous):

x+3y>-2 \[2x+y \ge 3\] so those are the two equations, now you need them in slope intercept form or y=mx+b. Okay?

OpenStudy (anonymous):

ok so do i gt y by itself

OpenStudy (anonymous):

Yup! now do that for both equations

OpenStudy (anonymous):

three y=x+-2

OpenStudy (anonymous):

okay so 3y=x+-2 Now you can divide both sides by 3 so you're left with y>1/3x+-2/3 For the first quation

OpenStudy (anonymous):

do i divide three into -2/three

OpenStudy (anonymous):

\[y > 1/3x -2/3\] -2/3 is a fraction so you don't need to divide it out

OpenStudy (anonymous):

well 1/three divide by three is = 1

OpenStudy (anonymous):

i hat this graphing it makes my head hurt but will keep trying